# Asymmetric Bias in Text-to-Image Generation with Adversarial Attacks

### 22 Dec 2023 v1

@@ 3.4 Proposed Attack Evaluation
To assess the success of a targeted entity swapping attack, we use a classifier to verify if the generated image matches the input or target prompt. Given a tuple (input text, target text, generated image), we define a classifier $\mathcal{C}$ as follows:
@@
We define class 00 as cases where the adversarial token search algorithm gets stuck at a local maximum, failing to reach the target within the compute budget. For example, when trying to change “A backpack in a forest” to “A cabin in a forest”, we noticed that some of the generated images depicted “People in a forest” or “A cabin and a backpack in a forest” instead. Class $+1$ alone indicates a successful attack, but this three-class framework enables a more comprehensive comparison between human judgments and our proposed classifiers.
Attack Success Rate (ASR).
Stable Diffusion supports the generation of multiple images in parallel. We define an adversarial suffix as being successful if the target text is a suitable caption for the majority of images (generated in parallel). For example, when trying to change “A backpack in a forest” to “A cabin in a forest”, if we generate 5 images with adversarial suffix appended prompt “A backpack in a forest titanic tycoon cottages caleb dojo, we will consider the adversarial suffix successful by majority vote if 3 or more images match the target prompt.
@@
We gather evaluations from three human evaluators for 200 random samples by presenting them a WebUI (Appendix E) with the generated image and two checkboxes for input text and target text. They are instructed to select texts that match the image and can select one, both, or neither, i.e. into three classes as established in Eqn. 4.
@@
Choice of the Classifier.
We generate multiple attack suffixes and images for each input-target pair to empirically determine attack success rates. Due to the large volume of images, we employ human evaluators for a subset and VLM-based classifiers for the full set evaluation. Specifically, we test InstructBLIP Liu et al. (2023a), LLaVA-1.5 Liu et al. (2023b), and CLIP Radford et al. (2021), and compare their performance with human labels.
@@
Figure 4: Comparison of pair-wise attack success rate on 50 pairs using Multiple Token Perturbation Algorithm.
For auto-regressive VLMs like InstructBLIP and LLaVA-1.5, we use the prompt ‘Does the image match the caption [PROMPT]? Yes or No?’, replacing [PROMPT] with input or target text for image evaluation. For CLIP models, an image is classed as $+1$ if its target text similarity is above $1-\gamma$ and its input text similarity is below $\gamma$, and $-1$ for the reverse case. All other cases are classified as 00.
Table 1 indicates that InstructBLIP outperforms other models such asLLaVA-1.5, CLIP-ViT-L14, and CLIP-ViT-L14-336. Given that CLIP models produce continuous values, we present their performance using the optimal threshold $\gamma$ ($\gamma_{CLIP}=0.0034$ and $\gamma_{CLIP-336}=0.0341$)that maximizes the F1 score against human labels. Consequently, we proceed with using InstructBLIP as the evaluation classifier in subsequent sections.
@@ 4 Experiments and Results
This section presents the experimental details and results of adversarial attacks for entity swapping setups, involving the insertion of adversarial suffixes.
@@ 4.1 Experimental Setups
We run all experiments, including the attacks, image generation, and VLM evaluation, on a single NVIDIA 3060 12GB and 16GB RAM. We use Stable Diffusion v2-1-base for all experiments unless stated otherwise.
We evaluate 100 input-target pairs to compare the effectiveness of Single and Multiple Token Perturbation. We run each algorithm 10 times per pair with $T=100$ steps with $k=5$ and $B=512$, which yields 10 adversarial attacks per pair. For each attack, we generate 5 images. The two algorithms are evaluated against each other on $100\times 10\times 5=5000$ generated images. We set $w_{t}=w_{s}=1$ in Eqn. 2 for the experiments.
@@ 4.2 Overall Attack Results
Using the same hyperparameters and compute budget, our Multiple Token Perturbation algorithm outperforms the Single Token Perturbation used in LLM attacks, with ASR $26.4\%$ and $24.4\%$ respectively over 100 input-target pairs on 5000 generated images. Zou et al. (2023) showed that Single Token Perturbation was an effective adversarial suffix-finding strategy for LLMs. However, the CLIP text is relatively lightweight compared to LLMs and behaves more like a bag-of-words model Yuksekgonul et al. (2022). We find that updating multiple tokens at each time step leads to faster convergence, likely because CLIP demonstrates a reduced emphasis on the semantic and syntactical relationships between tokens, in comparison to LLMs. Our findings corroborate the effectiveness of the Genetic Algorithm in Zhuang et al. (2023), which resembles multiple token perturbations but operates in an untargeted setting without a gradient-based algorithm. However, it’s important to note that Single Token Perturbation could potentially achieve the same ASR as Multiple Token Perturbation if not limited by compute resources. We employ Multiple Token Perturbation for all subsequent experiments analyzing asymmetric attacks.
@@ 4.3 Forward and Backward Attack Results
One of our key findings is the strong asymmetry of adversarial attack success rate, as illustrated in Figure 4. For instance, attacks from ‘A swan swimming in a lake.’ to ‘A horse swimming in a lake.’ failed in all ten attempts, whereas the reverse direction achieved an ASR of 0.9. Such asymmetric ASRs are common; however, in other cases, the forward and backward ASRs aren’t inversely proportional. For example, both directions between ‘A man reading a book in a library.’ and ‘A woman reading a book in a library.’ have moderate ASRs of 0.7 and 0.5, respectively, while pairs like (‘A dragon and a treasure chest.’, ‘A knight and a treasure chest.’) fail in both directions. Inspired by these asymmetric observations, we further conducted experiments to analyze the relationship between prompt distribution and attack success rate.
@@ 5 Asymmetric ASR Analysis
This section discusses our experiments to analyze the asymmetric ASR observed in Section 4.3. We aim to investigate the model’s internal beliefs that may lead to these distinct attack success rate (ASR) differences from opposite directions. We propose three potential factors for this asymmetry: the difficulty of generating the target text (BSR, Eqn. 5), the ‘naturalness’ of the target text relative to the input text ($\Delta_{1}$, Eqn. 6), and the difference in distance from the target text to the baseline compared to that from the input text ($\Delta_{2}$, Eqn. 7).
@@ 5.1 Probe Metrics
We initially speculated that the ASR might be related to the difficulty of generating the target prompt, leading us to evaluate the Base Success Rate (BSR) of target generation.
@@
BSR assesses T2I model’s likelihood of producing a target from a PAD token, for example, answering what would Stable Diffusion generate if prompted with ”A _____ swimming in a lake”. Figure 5 shows that T2I models have inherent biases towards certain objects, possibly due to biases in its training data. T2I models may struggle to generate images absent from their training data, as evidenced by our experiments where even simple scenes like “A dragon guarding a treasure” are inconsistently produced. Therefore, if T2I models struggle with the target alone, adversarial attacks aimed at generating them are likely to be even more challenging.
Figure 5: Base success rate: Images generated by Stable Diffusion (SD) 2.1 when prompted with the PAD token (”!” for SD 2.1).
(a) ASR vs. Perplexity Difference ($\Delta_{1}$ in Eqn. 6)
(b) ASR vs. BSR (of target text)
(c) ASR vs. Baseline Distance Difference ($\Delta_{2}$ in Eqn. 7)
(d) ASR for Negative and Positive $\Delta_{2}$
Figure 6: Correlation of ASR on $\Delta_{1}$, $\Delta_{2}$ and BSR. We find that the Perplexity Difference $\Delta_{1}$ does not correlate with ASR. BSR shows a weak positive correlation and Baseline Distance Difference $\Delta_{2}$ shows a moderate negative correlation with ASR.
We also speculated that the difference in Perplexity $\Delta_{1}$, measuring how natural or plausible a prompt is, might be associated with asymmetric ASR. For instance, “A swan swimming in a lake” is a more natural scene than “A horse swimming in a lake”. We selected the Perplexity (PPL) metric from text-davinci-003 by OpenAI Brown et al. (2020), as it most closely matches human judgment compared to T5 Raffel et al. (2020) and LLaMA 2 Touvron et al. (2023) in our experiments. Perplexity for a sequence $x_{1:n}$ is defined as $\text{PPL}(x_{1:n})=e^{-\frac{1}{n}\sum_{i=1}^{n}\log P(x_{i}|x_{1:i-1})}$, and we define the difference, noted as $\Delta_{1}$:
@@
Additionally, we introduce a new metric termed Baseline Distance Difference, denoted as $\Delta_{2}$. This combines insights from the probe metric BSR and $\Delta_{1}$, but measures distances using cosine similarity of CLIP embeddings instead of LLM’s perplexity.
@@ 5.2 Results
We evaluate 50 random prompt pairs in both directions, resulting in $50\times 2\times 10\times 5=5000$ generated images. To find the base success rate defined in Eqn. 5, we generate 64 images for each of the 50 random prompts and count the number of generations (out of 64) that were successful.
We find that $\Delta_{1}$ has a negligible correlation with ASR (Pearson $r=0.05$ and Spearman $\rho=-0.06$, see in Figure 5(a)). This is counter-intuitive because one might expect a target with lower perplexity compared to the input text should be easier to generate through an adversarial attack.
We also observed that ASR has a weak positive correlation with BSR (see Figure 5(b)) and a moderate one with $\Delta_{2}$ (see Figure 5(c)). In particular, Figure 5(d) shows the mean ASR when $\Delta_{2}<0$ is 0.40, whereas it drops to just 0.12 when $\Delta_{2}\geq 0$. Thus, $\Delta_{2}$ enables us to estimate, to some extent, the likelihood of a successful adversarial attack.
@@ 5.3 Predictor for Successful Attack
Considering the observed reasonable correlations of BSR (of the target text) and $\Delta_{2}$ with attack success rates, this section explores whether combining these two indicators can predict the likelihood of a successful entity-swapping attack.
BSR $\boldsymbol{\Delta_{2}}$ Samples Avg. ASR
$<0.9$ $<0$ 23 0.174
$<0.9$ $\geq 0$ 19 0.047
$\boldsymbol{\geq 0.9}$ $\boldsymbol{<0}$ 27 0.6
$\geq 0.9$ $\geq 0$ 31 0.172
Table 2: Average ASR for different combinations of BSR and $\Delta_{2}$. From the holdout set performance, we set $0.9$ as the threshold to classify BSR as high.
Table 2 demonstrates that our probe metric provides a reliable success predictor for attacks: when BSR (of the target text) is high and $\Delta_{2}$ is negative for a given input-target text pair, adversarial attacks have a 60% chance of success, dropping to about 5% when BSR is low and $\Delta_{2}$ is positive. Considering both BSR and $\Delta_{2}$ together improves the prediction of an attack’s success likelihood. However, the need for BSR precomputation is a limitation, and finding ways to approximate BSR without precomputation is an area for future research.
We also identified factors akin to existing research on general elements associated with attack success rates, like the length of the adversarial suffix. These factors, together with our experimental results, are detailed in Appendix D.
@@ 6 Conclusion
This paper presents an empirical study on adversarial attacks targeting text-to-image (T2I) generation models, with a specific focus on Stable Diffusion. We define a new attack objective—entity swapping—and introduce two gradient-based algorithms to implement the attack. Our research has identified key factors for successful attacks, revealing the asymmetric nature of attack success rates for forward and backward attacks in entity swapping. Furthermore, we propose probing metrics to associate the asymmetric attack success rate with the asymmetric bias within the T2I model’s internal beliefs, thus establishing a link between a model’s bias and its robustness against adversarial attacks.
Appendix A Algorithms
@@
Appendix B Changing the Number of Adversarial Tokens k
(a) Reducing the number of attack tokens ”a red panda/car in a forest.”. Only displaying the adversarial attack suffixes. 2 tokens are sufficient. ”a red panda in a forest. chained sportscar porsche generates ”a car in a forest”.
(b) Reducing the number of attack tokens ”a guitar/piano in a music store.”. Only displaying the adversarial attack suffixes. All 5 tokens are necessary. ”a guitar in a music store. serendipity upright three bank piano generates ”a piano in a music store.”
Figure 7: Reducing the number of tokens in adversarial prompts. Highly dependent on the starting-target text pair.
Appendix C T2I Model Basics
Stable Diffusion Rombach et al. (2022) is built on a denoising diffusion probabilistic model (DDPM) Ho et al. (2020) framework, utilizing a U-Net architecture for its core operations. Key to its text-to-image capabilities is the cross-attention mechanism, which aligns textual inputs with relevant visual features. Specifically, the U-Net attends to image-aligned text embeddings produced by a CLIP Radford et al. (2021) model. Stable Diffusion also incorporates a Variational Autoencoder Kingma and Welling (2013) to efficiently encode images into a latent space, significantly reducing computational requirements while maintaining image quality. Since text embedding generation using a CLIP model is the first stage in the Stable Diffusion pipeline, it is particularly susceptible to adversarial attacks (Galindo and Faria, ; Zhuang et al., 2023). If an adversary can perturb the text embeddings, later stages in the Stable Diffusion pipeline will reflect the perturbed embeddings.
C.1 Exploiting CLIP’s Embedding Space
The CLIP text-encoder maps the textual prompt tokens $x_{1:n}$, with $x_{i}\epsilon\{1,...,V\}$ where V denotes the vocabulary size, namely, the number of tokens to $x_{1:n}$, where $h_{i}$ is the hidden state corresponding to the token $x_{i}$. The U-Net component in Stable Diffusion attends to all $h_{1:n}$ embeddings using cross-attention. $x_{1:n}$ can be flattened into $\mathbb{\Phi}$, a one-dimensional vector of shape $n\times D$, where D is the embedding dimension (typically 768 for CLIP and its variants). For simplicity, we refer to $\mathbb{\Phi}$ as the text embedding of $x_{1:n}$ from here on. Let $\mathcal{H}$ represent the combined operation for encoding tokens $x_{1:n}$ and reshaping the output hidden states.
$\mathbb{\Phi}=\mathcal{H}(x_{1:n})=\text{Flatten}(\text{CLIP}(x_{1:n}))$ (8)

### 14 Feb 2024 v2

@@ 3.5 Proposed Attack Evaluation
To assess the success of a targeted entity-swapping attack, we use a classifier to verify if the generated image matches the input or target prompt. Given a tuple (input text, target text, generated image), we define a classifier $\mathcal{C}$ as follows:
@@
When trying to change “A backpack in a forest” to “A cabin in a forest”, we noticed that some of the generated images depicted “People in a forest” or “A cabin and a backpack in a forest” instead. We define such cases as class 00. Class $+1$ alone indicates a successful attack, but this three-class framework enables a more comprehensive comparison between human judgments and our proposed classifiers.
Attack Success Rate (ASR)
We define an adversarial suffix as successful if the target text is a suitable caption for the majority of images generated by an attack prompt using a T2I model. For example, if we generate 5 images with an appended adversarial suffix prompt “A backpack in a forest.titanic tycoon cottages caleb dojo, we will consider the adversarial suffix successful if 3 or more images match the target prompt “A cabin in a forest”.
@@
Choice of the Classifier
We generate multiple attack suffixes for each input-target pair to determine attack success rates. Due to the large volume of images, we employ human evaluators for a subset and VLM-based classifiers for the full set evaluation. We test InstructBLIP Liu et al. (2023a), LLaVA-1.5 Liu et al. (2023b) and CLIP Radford et al. (2021), and compare their performance with human labels.
@@
Figure 4: Comparison of pair-wise attack success rate on HQ-Pairs using Multiple Token Perturbation Algorithm.
For InstructBLIP and LLaVA-1.5, we use the prompt ‘Does the image match the caption [PROMPT]? Yes or No?’. For CLIP models, an image is classified as $+1$ if its target text similarity is above $1-\gamma$ and its input text similarity is below $\gamma$ and $-1$ for the reverse case. All other cases are classified as 00. Table 1 shows the agreement of different automatic classifiers with ground truths from our human evaluators. We use the optimal threshold $\gamma$ ($\gamma_{CLIP}=0.0034$ and $\gamma_{CLIP-336}=0.0341$) that maximizes the F1 score. Since InstructBLIP shows the best alignment with human evaluation, we use InstructBLIP as our sole classifier in subsequent sections.
@@ 4 Experiments and Results
This section presents the experimental details and results of adversarial attacks for entity-swapping, involving the insertion of adversarial suffixes.
@@ 4.1 Experimental Setups
We evaluate Stable Diffusion v2-1-base on the HQ-Pairs dataset of 100 input-target pairs to compare the effectiveness of Single and Multiple Token Perturbation. We run each algorithm 10 times per pair with $T=100$ steps with $k=5$ and $B=512$, which yields 10 adversarial attacks per pair, and we generate 5 images per attack. The two algorithms are evaluated against each other on $100\times 10\times 5=5000$ generated images. We set $w_{t}=w_{s}=1$ in Eqn. 2 for the experiments. Afterward, we evaluate COCO-Pairs (1000 pairs) using the Multiple Token Perturbation algorithm to establish the asymmetric bias phenomenon with the same hyperparameters. We used a single Nvidia RTX 4090 GPU for all experiments, including attack, image generation, and automated evaluation, totaling around 500 GPU hours.
@@ 4.2 Overall Attack Results
Using the same hyperparameters and compute budget, our Multiple Token Perturbation algorithm outperforms the Single Token Perturbation ( ASR $26.4\%$ vs. $24.4\%$ for 1000 attacks). Zou et al. (2023) showed that Single Token Perturbation was an effective adversarial suffix-finding strategy for LLMs. However, the CLIP text is relatively lightweight compared to LLMs and behaves more like a bag-of-words model Yuksekgonul et al. (2022). CLIP also has a larger vocabulary compared to LLMs (${}50K$ vs. $32K$) which leads to a larger unrestricted search space ($\sim 10^{24}$ vs. $\sim 10^{23}$ for 5 token suffixes). We find that updating multiple tokens at each time step leads to faster convergence, likely because CLIP demonstrates a reduced emphasis on the semantic and syntactical relationships between tokens. Our findings corroborate the effectiveness of the Genetic Algorithm in Zhuang et al. (2023), which resembles multiple token perturbations but operates in an untargeted setting without a gradient-based algorithm. We employ Multiple Token Perturbation for all subsequent experiments.
@@ 4.3 Forward and Backward Attack Results
One of our key findings is the strong asymmetry of adversarial attack success rate, as illustrated in Figure 4. For instance, attacks from ‘A swan swimming in a lake.’ to ‘A horse swimming in a lake.’ failed in all ten attempts, whereas the reverse direction achieved an ASR of 0.9. In other cases, the forward and backward ASRs aren’t inversely proportional. For example, both directions between ‘A man reading a book in a library.’ and ‘A woman reading a book in a library.’ have moderate ASRs of 0.7 and 0.5, respectively, while pairs like (‘A dragon and a treasure chest.’, ‘A knight and a treasure chest.’) fail in both directions. Inspired by these asymmetric observations, we conducted further experiments to analyze the relationship between prompt distribution and attack success rate.
@@ 5 Asymmetric ASR Analysis
This section discusses our experiments to analyze the asymmetric ASR observed in Section 4.3. We aim to investigate the model’s internal beliefs that may lead to these distinct attack success rate (ASR) differences from opposite directions. We propose three potential factors for this asymmetry: the difficulty of generating the target text (BSR, Eqn. 5), the naturalness of the target text relative to the input text ($\Delta_{1}$, Eqn. 6), and the difference in distance from the target text to the baseline compared to that from the input text ($\Delta_{2}$, Eqn. 7).
@@ 5.1 Probe Metrics
We initially speculated that ASR might be related to the difficulty in generating the target prompt, leading us to evaluate the Base Success Rate (BSR) of target generation.
@@
BSR assesses the T2I model’s ability to generate an image that matches the input prompt without any adversarial suffixes. Stable Diffusion is often unable to generate novel compositions not present in its training data West et al. (2023) and struggles with generating co-hyponym entities in the same scene Tang et al. (2022). We find that even simple scenes such as “A dragon guarding a treasure.” are inconsistently produced (See Appendix F for examples). Therefore, if the T2I models struggle with the target alone, adversarial attacks aimed at generating them are likely to be even more challenging.
Figure 5: Baseline Distance Difference measures the inherent biases of T2I models. This can be observed by prompting Stable Diffusion a PAD token in place of an entity.
(a) ASR vs. Baseline Distance Difference ($\Delta_{2}$ in Eqn. 7)
(b) ASR for Negative and Positive $\Delta_{2}$
Figure 6: Correlation of ASR with Baseline Distance Difference $\Delta_{2}$. Data is reported using the Multiple Token Perturbation algorithm on HQ-Pairs. $\Delta_{2}$ shows a moderate negative correlation with ASR.
We also speculated that the difference in Perplexity $\Delta_{1}$, measuring how natural or plausible a prompt is, might be associated with asymmetric ASR. For example, “A swan swimming in a lake” is a more natural scene than “A horse swimming in a lake”. Using text-davinci-003 by OpenAI Brown et al. (2020), we calculate the perplexity difference
@@
where $\text{PPL}(x_{1:n})=e^{-\frac{1}{n}\sum_{i=1}^{n}\log P(x_{i}|x_{1:i-1})}$ is the perplexity for the sequence $x_{1:n}$.
Furthermore, we introduce a new metric termed Baseline Distance Difference, denoted as $\Delta_{2}$. Figure 5 shows that T2I models have inherent biases towards certain objects. We denote this phenomenon as the baseline - answering what would Stable Diffusion generate if prompted with “A _____ swimming in a lake”. Intuitively, targets closer to this baseline should be easier to generate.
@@ 5.2 Results
We generated 64 images for each sentence in HQ-Pairs and COCO-Pairs. We counted the number of successful generations to determine the BSR as defined in Eqn. 5.
On the HQ-Pairs dataset, we find that Perplexity Difference $\Delta_{1}$ has a negligible correlation with ASR (Pearson $r=0.05$ and Spearman $\rho=-0.06$). This is counterintuitive because we expected that a target with lower perplexity compared to the input text would be easier to generate through an adversarial attack. We also observed that ASR has a weak positive correlation with BSR (Pearson $r=0.28$ and Spearman $\rho=0.38$) and a moderate correlation with $\Delta_{2}$ (Pearson $r=-0.39$ and Spearman $\rho=-0.46$. See Figure 5(a)). In particular, Figure 5(b) shows that the mean ASR is 0.40 when $\Delta_{2}$ is negative, while it drops to just 0.12 when $\Delta_{2}$ is positive. Thus, $\Delta_{2}$ allows us to estimate, to some extent, the probability of a successful adversarial attack. We present more correlation plots of ASR with Perplexity Difference and BSR in Appendix F.
@@ 5.3 Predictor for Successful Attack
Considering the observed correlations of BSR (of the target text) and $\Delta_{2}$ with attack success rates, this section explores whether the combination of these two indicators can predict the probability of a successful entity-swapping attack.
HQ-Pairs COCO-Pairs
BSR $\Delta_{2}$ Num. Avg. ASR Num. Avg. ASR
Low Neg. 23 0.174 260 0.129
Low Pos. 19 0.047 274 0.087
High Neg. 27 0.6 239 0.349
High Pos. 31 0.171 226 0.213
All All 100 0.264 1000 0.189
Table 2: Average ASR for different combinations of BSR and $\Delta_{2}$ on COCO-Pairs dataset. We define $\text{BSR}\geq 0.9$ as high. The average BSR of the target text of HQ-Pairs and COCO-Pairs were 0.82 and 0.698 respectively.
Table 2 shows that our probe metric acts as a reliable predictor of attack success: when BSR (of the target text) is high and $\Delta_{2}$ is negative for a given input-target text pair, adversarial attacks have a 60% chance of success on the HQ-Pairs dataset, compared to only 5% when BSR is low and $\Delta_{2}$ is positive. Thus, considering both BSR and $\Delta_{2}$ together enhances the prediction accuracy of an attack’s success likelihood. We further validate our findings on the much larger COCO-Pairs dataset. Although the differences are not as pronounced as those in the HQ-Pairs, due to limitations explained in Section 3.2, we still observe that high BSR and negative $\Delta_{2}$ remain indicative of a higher likelihood of successful adversarial attacks. We also identified factors akin to existing research on general elements associated with attack success rates, like the length of the adversarial suffix. These factors, together with our experimental results, are detailed in Appendix G.
@@ 6 Conclusion
This paper presents an empirical study on adversarial attacks targeting text-to-image (T2I) generation models, with a specific focus on Stable Diffusion. We define a new attack objective: entity-swapping, and introduce two gradient-based algorithms to implement the attack. Our research has identified key factors for successful attacks, revealing the asymmetric nature of attack success rates for forward and backward attacks in entity-swapping. Furthermore, we propose probing metrics to associate the asymmetric attack success rate with the asymmetric bias within the T2I model’s internal beliefs, thus establishing a link between a model’s bias and its robustness against adversarial attacks.
7 Limitations
[…]
One of our key findings is that asymmetric bias is not intuitive. Although humans might consider “fish” to be a more natural option (and likely more abundant in the training data) for “A ____ in an aquarium”, we find that Stable Diffusion is strongly biased towards “turtle” instead. We leave exploring the underlying reason for this non-intuitive bias as future work.
We observed that gradient-based algorithms tend to include the target word in the adversarial suffix. Concurrent works that aim to generate undetectable NSFW attacks use a dictionary to prevent this. Since we target benign words and have different targets for every attack, we could not use a similar approach. We explore explicitly forbidding tokens corresponding to the target word, but the algorithm still finds synonyms or different tokenizations of the target word. Forbidding the target word proved to be a nontrivial and ultimately, we did not consider generating true adversarial attacks to be a central focus of our investigation of model bias. Another technical challenge is the need to compute BSR which involves generating a statistically significant number of images (64 in our experiments) for the same prompt. Finding ways to approximate the BSR is an area for future research.
Appendix A Generating COCO-Pairs
Starting from 5000 captions, we filter out long captions and use a Named-Entity-Recognition model Nadeau and Sekine (2007) to identify the first noun in the sentence and use a Fill-Mask model Devlin et al. (2018) to replace it with another noun. We use the NLTK Loper and Bird (2002) library and several heuristics to prevent synonyms, hyponym-hypernym, and nonvisualizable nouns from being selected. We are left with 2093 (base caption, synthetic caption) pairs, from which we sample 500. This yields 1000 sentence pairs in total by considering both directions.
Appendix B Algorithms
@@
Appendix C Additional Examples of Asymmetric Bias
Sentence Pair (1 / 2) $\boldsymbol{\Delta_{2}}$
ASR $\mathbf{{\color[rgb]{0,0,1}1}\rightarrow{\color[rgb]{1,.5,0}2}}$
ASR $\mathbf{{\color[rgb]{1,.5,0}2}\rightarrow{\color[rgb]{0,0,1}1}}$ Example
a (plane / hot air balloon) in the sky at sunset. -0.1 80% 0% [Uncaptioned image]
a (cabin / backpack) on a mountain. -0.08 90% 20% [Uncaptioned image]
an owl in a (forest / shopping mall). -0.07 70% 0% [Uncaptioned image]
a (birdhouse / jack o lantern) on a tree branch. -0.06 60% 0% [Uncaptioned image]
a (turtle / fish) swimming in an aquarium. +0.05
0%
90% [Uncaptioned image]
a (robot / human) dancing in the rain. +0.1
0%
60% [Uncaptioned image]
a doodle of a (light bulb / dog) on a blackboard. +0.1
0%
80% [Uncaptioned image]
Table 3: Additional examples of asymmetric bias in Stable Diffusion 2.1. $\Delta_{2}$ shows a consistent negative correlation with ASR.
Appendix D Changing the Number of Adversarial Tokens
(a) Reducing the number of attack tokens for ”a red panda/car in a forest.”. Displaying only the adversarial attack suffixes. 2 tokens are sufficient. ”a red panda in a forest.chained porsche generates ”a car in a forest”.
(b) Reducing the number of attack tokens ”a guitar/piano in a music store.”. Displaying only the adversarial attack suffixes. All 5 tokens are necessary. ”a guitar in a music store. serendipity upright three bank piano generates ”a piano in a music store.”
Figure 7: Reducing the number of tokens in adversarial prompts. Highly dependent on the input-target text pair.
Appendix E T2I Model Basics
Stable Diffusion Rombach et al. (2022) is built on a denoising diffusion probabilistic model (DDPM) Ho et al. (2020) framework, utilizing a U-Net architecture for its core operations. Key to its text-to-image capabilities is the cross-attention mechanism, which aligns textual inputs with relevant visual features. Specifically, the U-Net attends to image-aligned text embeddings produced by a CLIP Radford et al. (2021) model. Stable Diffusion also incorporates a Variational Autoencoder Kingma and Welling (2013) to efficiently encode images into a latent space, significantly reducing computational requirements while maintaining image quality. Since text embedding generation using a CLIP model is the first stage of the Stable Diffusion pipeline, it is particularly susceptible to adversarial attacks (Galindo and Faria, ; Zhuang et al., 2023). If an adversary can perturb the text embeddings, later stages in the Stable Diffusion pipeline will reflect the perturbed embeddings.
E.1 Exploiting CLIP’s Embedding Space
The CLIP text-encoder maps the textual prompt tokens $x_{1:n}$, with $x_{i}\epsilon\{1,...,V\}$ where V denotes the vocabulary size, namely, the number of tokens to $x_{1:n}$, where $h_{i}$ is the hidden state corresponding to the token $x_{i}$. The U-Net component in Stable Diffusion attends to all $h_{1:n}$ embeddings using cross-attention. $x_{1:n}$ can be flattened into $\mathbf{\Phi}$, a one-dimensional vector of shape $n\times D$, where D is the embedding dimension (typically 768 for CLIP and its variants). For simplicity, we refer to $\mathbf{\Phi}$ as the text embedding of $x_{1:n}$ from here on. Let $\mathcal{H}$ represent the combined operation for encoding tokens $x_{1:n}$ and reshaping the hidden output states.
$\mathbf{\Phi}=\mathcal{H}(x_{1:n})=\text{Flatten}(\text{CLIP}(x_{1:n}))$ (8)